OpenStudy (anonymous):

i need help solving quadratic equations. (k+1)(k-5)=0

OpenStudy (mertsj):

If you multiply two numbers and get 0 then one of the numbers has to be 0. You are multiplying the two numbers k+1 and k-5. One or both of them has to be zero. So set each one to 0 and solve the resulting equations.

OpenStudy (anonymous):

can you walk me through it i don't understand.

OpenStudy (anonymous):

what Mertsj is saying : (k+1)(k-5)=0 => (k+1)=0 or (k-5)=0 k+1=0 => k=-1 k-5=0 => k=5